OpenStudy (anonymous):

Check my answer plz!Ill medal! A miner descended 1200 feet into a mine. While in the mine, he then rose 900 feet and then descended 350 feet. What was the miner's net gain or loss in altitude? A. –650 ft B. 50 ft C. –1200 ft D. 650 ft (I pick D)

Vocaloid (vocaloid):

not quite, check your sign

OpenStudy (anonymous):

A then

OpenStudy (anonymous):

-1200 +900 -350 = ?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

It's best to draw/record/write down everything you see in a word problem that may come into use. That's what I did to demonstrate this problem. What do you think it is again?

OpenStudy (anonymous):

Incorrect. What Vocaloid meant was to check your signs in the initial application of this problem.

OpenStudy (anonymous):

i have no clue

OpenStudy (anonymous):

Start off with -1200 + 900 = ?

OpenStudy (anonymous):

You can even apply it into a calculator w/o the signs… just make sure to keep the negative sign if the greater number is a negative

OpenStudy (anonymous):

-300

OpenStudy (anonymous):

Good job! To make it easier-looking with less complex numbers, what's -5 + 3?

OpenStudy (anonymous):

The same thing is applied to any number :)

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

Awesome! So -300 - 350 = ? I think you may have mixed up the numbers with multiplication because a negative times a negative = positive. So it equals to -650, which is A. I apologize for saying it was wrong.. that was my error.

OpenStudy (anonymous):

So awesome job! A is the correct answer :)

OpenStudy (anonymous):

Tag me if you need anymore help! :)

OpenStudy (anonymous):

ty